a=b
ab=b2
ab-a2 = b2-a2
a(b-a)=(b+a)(b-a)
a=b+a
a=a+a
a=2a
1=2
My teacher says that 1=2 but this problem is NOT correct, can you explain to my why it will not work?

Answer by jim_thompson5910(35256)

Since you're assuming that a=b, this means that b-a=a-a=0. So when you divide both sides by b-a, you're basically dividing by zero (which is NOT allowed).


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider a making a donation at my website. Thanks.

Jim